Grade 12Algebra

Find the angles which are equal to its complement??????

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let the measure of the required angle be xo.
We know that, sum of measures of complementary angle pair is 90o.
Then,
= x + x = 90o
= 2x = 90o
= x = 90/2
= x = 45o
Hence, the required angle measures is 45o.

Thanks